# Diffusion Posterior Proximal Sampling for Image Restoration

### v1

@@ 5.3 The Effect of Proximal Sampling
More Robust to Hyper-parameter.  We anticipate that our proposed proximal sampling serves as an adaptive correction for prediction errors resulting from imprecise parameter settings. To substantiate this claim, we conducted experiments with different $\lambda_{t}$, comparing the results of our proximal sampling with those obtained through random sampling.
Figure 6: FID comparison between random sampling and proximal sampling with different values of $\lambda_{t}$.
As illustrated in Figure 6, our method consistently outperforms random sampling across different parameter settings. Notably, the performance of the random sampling method exhibits fluctuations with changes in parameters. In contrast, our method demonstrates stability over a range of parameter configurations, thereby validating the assertion.
5.4 Ablation Study
The Number of Candidate Samples.  The number of samples, denoted as $n$, stands out as a pivotal parameter of the proposed methodology. The ablation studies relating to $n$ are shown in Table 2 and Figure 5. Figure 5 illustrates a noticeable enhancement in output quality and details with increasing $n$. This observation is reinforced by Table 2, where both the LPIPS and FID metrics exhibit a notable decrease as $n$ increases. The augmented number of candidate samples corresponds to a higher likelihood of proximity to the intended target, thereby providing enhanced supervision for the generative process. This reaffirms the reliability and applicability of the proposed methodology. It is noteworthy that the improvement becomes relatively marginal when $n$ exceeds 20 and may even result in a decrease in PSNR. We posit that measurement noise ($\mathbf{n}$ in Eq. (5)) influences the precision of our approximation. Moreover, sample selection is conducted within the range space, introducing a certain degree of misalignment with the null space. Consequently, we set $n=20$ in our experiments to strike a balance between perceptual metrics, distortion metrics, and speed.
@@ 5.4 Ablation Study
Table 2: Ablation studies on the number of candidate samples.
@@
$n$$*{}^{*}$ 26.93 0.203 26.30 24.79 0.276 33.75
$n$=100 26.83 0.201 25.79 24.72 0.268 32.91
Noise Level.  The precision of the approximation is impacted by the level of measurement noise $\mathbf{n}$. We explore the effects of different noise levels on the proposed method compared with random sampling. The results presented in Table 3 reveal that (i) both methods experience a decline in performance as the noise level increases, and our approach consistently outperforms random sampling in all settings. (ii) For inpainting, both methods experience considerable degradations as the noise level increases, demonstrating sensitivity to the noise levels. (iii) For SR, notable performance degradation is observed in our method as the noise level increases, while the method adopting random sampling demonstrates greater robustness in this regard.
Design Components.  The proposed method consists of two design components: aligned initialization and proximal sample selection. To evaluate the impact of these components on overall performance, we conducted ablation experiments on the inpainting task. The results, as presented in Table 4, highlight that the proximal sample selection is a pivotal component, significantly contributing to performance improvements. While our aligned initialization method contributes moderately to the enhancement, more analyses can be found in Section C.1.
Table 3: Comparison of the impact of different noise levels.
@@
6 Conclusion
In this paper, we introduce a novel approach to mitigate problems caused by misaligned random sampling in diffusion-based image restoration methods. Specifically, our approach advocates selecting the proximal sample that is more consistent with the measurement in the predicted distribution at each generative step. In addition, we propose a realignment of the inference initialization involving measurement information to better align with expected generations. Experimental results validate a substantial performance improvement compared to SOTA methods. Our method innovatively takes advantage of stochastic sampling in the diffusion generative process to enhance the generation quality, offering new insights for future work.
7 Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Score-based Diffusion Models
$\mathrm{d}\mathbf{x}=\mathbf{f}(\mathbf{x},t)\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},$ (17)
@@
$\mathrm{d}\mathbf{x}=[\mathbf{f}(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})]\mathrm{d}t+g(t)\mathrm{d}\mathbf{\bar{w}},$ (18)
@@
$\min_{\mathbf{\theta}}\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t}}\big{[}\|\mathbf{s}_{\mathbf{\theta}}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})\|_{2}^{2}\big{]},$ (19)
@@ B.1 Derivation for $\mathbf{x}_{t-1}^{*}$
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}^{*})={\cal N}(\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0}^{*}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\frac{\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}^{*}}{\sqrt{1-\bar{\alpha}_{t}}},\sigma_{t}^{2}\mathbf{I}).$ (20)
@@
$\mathbf{x}_{t-1}^{*}=$ $\sqrt{\bar{\alpha}_{t-1}}{\mathbf{x}}_{0}^{*}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}^{*}}{\sqrt{1-\bar{\alpha}_{t}}}$ (21)
$=$ $\frac{\sqrt{1-\bar{\alpha}_{t-1}}}{\sqrt{1-\bar{\alpha}_{t}}}\mathbf{x}_{t}+(\sqrt{\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t}}\cdot\sqrt{1-\bar{\alpha}_{t-1}}}{\sqrt{1-\bar{\alpha}_{t}}})\mathbf{x}_{0}^{*}.$
@@
$\mathcal{A}\mathbf{x}_{t-1}^{*}=\frac{\sqrt{1-\bar{\alpha}_{t-1}}}{\sqrt{1-\bar{\alpha}_{t}}}\mathcal{A}\mathbf{x}_{t}+(\sqrt{\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t}}\cdot\sqrt{1-\bar{\alpha}_{t-1}}}{\sqrt{1-\bar{\alpha}_{t}}})\mathcal{A}\mathbf{x}_{0}^{*}.$ (22)
By applying $\mathbf{y}=\mathcal{A}\mathbf{x}_{0}^{*}+\mathbf{n}$,
$\mathcal{A}\mathbf{x}_{t-1}^{*}=$ $\frac{\sqrt{1-\bar{\alpha}_{t-1}}}{\sqrt{1-\bar{\alpha}_{t}}}\mathcal{A}\mathbf{x}_{t}+(\sqrt{\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t}}\cdot\sqrt{1-\bar{\alpha}_{t-1}}}{\sqrt{1-\bar{\alpha}_{t}}})(\mathbf{y}-\mathbf{n})$ (23)
$\approx$ $\frac{\sqrt{1-\bar{\alpha}_{t-1}}}{\sqrt{1-\bar{\alpha}_{t}}}\mathcal{A}\mathbf{x}_{t}+(\sqrt{\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t}}\cdot\sqrt{1-\bar{\alpha}_{t-1}}}{\sqrt{1-\bar{\alpha}_{t}}})\mathbf{y}$
@@
Assume that $\|\mathbf{y}\|_{2}\leq Y,\|\mathbf{x}_{0}^{*}\|_{2}\leq X,\|\mathbf{x}_{t-1}^{\prime}\|_{2}\leq X,\|\bm{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})\|_{2}\leq E$ are bounded and continuous, there exists a upper bound for $\|\mathcal{A}\left(\mathbf{x}_{t-1}^{i}-\mathbf{x}^{*}_{t-1}\right)\|_{2}^{2}$ which is:
$\|\mathcal{A}\left(\mathbf{x}_{t-1}^{i}-\mathbf{x}^{*}_{t-1}\right)\|_{2}^{2}\leq\left(\mathcal{A}E-\mathcal{A}X+\mathcal{A}\sigma_{t}\right)^{2}.$ (24)
@@
$\|\mathcal{A}\left(\mathbf{x}_{t-1}^{i}-\mathbf{x}^{*}_{t-1}\right)\|_{2}^{2}$ $=\|\mathcal{A}\mathbf{x}_{t-1}^{i}-\mathcal{A}\mathbf{x}^{*}_{t-1}\|_{2}^{2}$ (25)
$\overset{(\text{a})}{\approx}\|\mathcal{A}\left(\bm{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}^{i}_{t}\right)-\left(\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\right)\|_{2}^{2},$ (26)
@@
$\|\mathcal{A}\left(\mathbf{x}_{t-1}^{i}-\mathbf{x}^{*}_{t-1}\right)\|_{2}^{2}$ $\overset{(\text{a})}{\approx}\|\mathcal{A}\left(\bm{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}^{i}_{t}\right)-\left(\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\right)\|_{2}^{2}$ (27)
$\approx\|\left(\mathcal{A}\bm{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})-\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}\right)+\left(\sigma_{t}\mathcal{A}\mathbf{z}_{t}^{i}-\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\right)\|_{2}^{2}$ (28)
$\approx\left(\|\mathcal{A}\bm{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})-\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}\|_{2}+\|\sigma_{t}\mathcal{A}\mathbf{z}_{t}^{i}-\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\|_{2}\right)^{2}.$ (29)
Since $\|\bm{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})\|_{2}\leq E$, $\|\bar{\alpha}_{t}\|_{2}\leq A$, $\|\bar{\alpha}_{t}\|_{2}\leq A$, $\|\bar{\alpha}_{t-1}\|_{2}\leq A,$, and $\|\mathbf{y}\|_{2}\leq Y$, we have the upper bound for $\|\mathcal{A}\bm{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})-\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}\|_{2}$, and thus $\|\mathcal{A}\bm{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})-\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}\|_{2}\leq\mathcal{A}E-\sqrt{A}Y$. Similarly, we have the upper bound for $\|\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\|_{2}$, which is $\|\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\|_{2}\leq\mathcal{A}X-\sqrt{A}Y$. Therefore, we have:
$\|\mathcal{A}\left(\mathbf{x}_{t-1}^{i}-\mathbf{x}^{*}_{t-1}\right)\|_{2}^{2}$ $\overset{(\text{a})}{\approx}\|\mathcal{A}\left(\bm{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}^{i}_{t}\right)-\left(\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\right)\|_{2}^{2}$ (30)
$\leq\left(\mathcal{A}E-\mathcal{A}X+\mathcal{A}\sigma_{t}\right)^{2}.$ (31)
@@
$f(\mathbf{z})=\|\mathcal{A}(\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}-C_{1}\mathbf{x}_{t})-C_{2}\mathbf{y}\|^{2}_{2},$ (32)
[…]
@@
$\hat{\mu}_{\text{MC}}=\frac{1}{N}\sum^{N}_{i=1}f(\mathbf{z}_{i}).$ (34)
@@
$\hat{\mu}_{\text{Ours}}=\operatorname*{arg\,min}_{\mathbf{z}}{f(\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{N})}.$ (35)
@@
$\text{Var}_{\text{DPS}}\left(f(\mathbf{z})\right)>\text{Var}_{\text{MC}}\left(f(\mathbf{z})\right)=\frac{\text{Var}_{\text{DPS}}\left(f(\mathbf{z})\right)}{N}>\text{Var}_{\text{Ours}}\left(f(\mathbf{z})\right).$ (36)
Appendix C Additional Ablation Studies
C.1 Design Components.
The proposed method consists of two design components: aligned initialization and proximal sample selection. To assess the impact of these components on overall performance, we conducted ablation experiments on the inpainting task. The results are presented in Table 4. Our observations are as follows: (i) compared with random initialization, the improvement from our initialization method is limited. We attribute this to the high NFEs of our method, making initialization less critical. (ii) Sample selection emerges as the key component, facilitating substantial performance enhancements. (iii) The best results are achieved when both components are combined. Consequently, minimizing randomness and fluctuations proves highly beneficial for restoration performance.
Table 4: Ablation studies on the design components of inpainting with $\sigma_{y}=0.01$.
Components FFHQ ImageNet

### v2

@@ 5.3. The Effect of Proximal Sampling
We anticipate that our proposed proximal sampling serves as an adaptive correction for prediction errors resulting from imprecise parameter settings. To substantiate this claim, we conducted experiments with different $\lambda_{t}$, comparing the results of our proximal sampling with those obtained through random sampling. As illustrated in Figure 6, our method consistently outperforms random sampling across different parameter settings. Notably, the performance of the random sampling method exhibits fluctuations with changes in parameters. In contrast, our method demonstrates stability over a range of parameter configurations, thereby validating the assertion.
5.4. Computational Resource
To examine the impact of the variable $n$ on computational efficiency, we conducted an evaluation that consisted of measuring memory consumption and computational time required for the SR×$\times 4$ task. This evaluation was performed on a single NVIDIA RTX 3090 GPU, and the results are presented in Table 2. Notably, when the value of $n$ is relatively small ($\leq$20), our method shows a substantial enhancement in the restoration metric as $n$ increases while incurring minimal increases in memory consumption and computation time. Additionally, the carefully designed adaptive sampling frequency strategy further enhances restoration efficiency within less computation time. These findings highlight the effectiveness and validity of our approach.
@@ 5.4. Computational Resource
Table 3. Ablation studies on the number of candidate samples on FFHQ SR×$\times$ 4.
@@
$n$=20 26.93 0.203 26.30 24.79 0.276 33.75
adaptive $n$ 26.94 0.201 25.98 24.44 0.267 30.70
5.5. Ablation Study
5.5.1. The Number of Candidate Samples.
The number of samples, denoted as $n$, stands out as a pivotal parameter of the proposed methodology. The ablation studies relating to $n$ are shown in Table 3 and Figure 5. Figure 5 illustrates a noticeable enhancement in output quality and details with increasing $n$. This observation is reinforced by Table 3, where both the LPIPS and FID metrics exhibit a notable decrease as $n$ increases. The augmented number of candidate samples corresponds to a higher likelihood of proximity to the intended target, thereby providing enhanced supervision for the generative process. The enhancement of the proposed adaptive sampling frequency further enhance the efficiency and applicability of our proposed method.
It is noteworthy that the improvement becomes relatively marginal when $n$ exceeds 20 and may even lead to a decrease in PSNR. There are several reasons for this phenomenon. (i) The approximation is only conducted within the range space, neglecting the null space. This can lead to a certain degree of misalignment with the reference image. (ii) Since we utilized $\mathbf{y}\approx\mathcal{A}\mathbf{x}_{0}$ in our algorithm, the presence of measurement noise ($\mathbf{n}$ in Eq. (5)) affects the accuracy of our sample selection. (iii) It can also be attributed to the well-known trade-off between perception and distortion metrics (Blau and Michaeli, 2018). Consequently, a larger value of $n$ does not necessarily lead to better results.
5.5.2. Noise Level.
The precision of the approximation is impacted by the level of measurement noise $\mathbf{n}$. We explore the effects of different noise levels on the proposed method compared with random sampling. The results presented in Table 4 reveal that (i) both methods experience a decline in performance as the noise level increases, and our approach consistently outperforms random sampling in all settings. (ii) For inpainting, both methods experience considerable degradations as the noise level increases, demonstrating sensitivity to the noise levels. (iii) For SR, notable performance degradation is observed in our method as the noise level increases, while the method adopting random sampling demonstrates greater robustness in this regard.
Table 4. Comparison of the impact of different noise levels.
@@ 5.5. Ablation Study
6. Conclusion
In this paper, we introduce a novel approach to mitigate problems caused by misaligned random sampling in diffusion-based image restoration methods. Specifically, our approach advocates selecting the proximal sample that is more consistent with the observed measurement in the predicted distribution. An adaptive sampling frequency strategy is followed to optimize the computational efficiency of the proposed method. In addition, we propose a realignment of the inference initialization involving measurement information to better align with expected generations. Experimental results validate a substantial performance improvement compared to SOTA methods. Our method innovatively takes advantage of stochastic sampling in the diffusion generative process, exploiting the sampling selection to enhance the generation quality and offering new insights for future diffusion inference algorithms.
Acknowledgements.
This work was supported by the Fundamental Research Funds for the Central Universities (No. 1082204112364), the National Nature Science Foundation of China (No. 62106161), and the Sichuan University Luzhou Municipal Government Strategic Cooperation Project (No. 2022CDLZ-8). We thank the anonymous reviewers for their insightful comments and feedback, and extend our gratitude to Mengxi Xie from SJTU and Linrui Dai from CQU for their constructive suggestions on the figures.
@@ Appendix A Score-based Diffusion Models
(18) $\mathrm{d}\mathbf{x}=\mathbf{f}(\mathbf{x},t)\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},$
@@
(19) $\mathrm{d}\mathbf{x}=[\mathbf{f}(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})]\mathrm{d}t+g(t)\mathrm{d}\mathbf{\bar{w}},$
@@
(20) $\min_{\mathbf{\theta}}\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t}}\big{[}\|\mathbf{s}_{\mathbf{\theta}}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})\|_{2}^{2}\big{]},$
@@ B.1. Derivation for $\mathbf{x}_{t-1}^{*}$
(21) $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$
[…]
@@
(22) $\mathbf{x}_{t-1}^{*}={}$ $\sqrt{\bar{\alpha}_{t-1}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}}{\sqrt{1-\bar{\alpha}_{t}}}$
[…]
@@
(23) $\mathcal{A}\mathbf{x}_{t-1}^{*}=\frac{\sqrt{1-\bar{\alpha}_{t-1}}}{\sqrt{1-\bar{\alpha}_{t}}}\mathcal{A}\mathbf{x}_{t}+(\sqrt{\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t}}\cdot\sqrt{1-\bar{\alpha}_{t-1}}}{\sqrt{1-\bar{\alpha}_{t}}})\mathcal{A}\mathbf{x}_{0}.$
By applying $\mathbf{y}=\mathcal{A}\mathbf{x}_{0}+\mathbf{n}$,
(24) $\mathcal{A}\mathbf{x}_{t-1}^{*}$
[…]
@@
Assume that $\|\mathbf{y}\|_{2}\leq Y,\|\mathbf{x}_{0}\|_{2}\leq X,\|\mathbf{x}_{t-1}^{\prime}\|_{2}\leq X,\|\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})\|_{2}\leq E$ are bounded and continuous, there exists a upper bound for $\|\mathcal{A}\left(\mathbf{x}_{t-1}^{i}-\mathbf{x}^{*}_{t-1}\right)\|_{2}^{2}$ which is:
(25) $\|\mathcal{A}\left(\mathbf{x}_{t-1}^{i}-\mathbf{x}^{*}_{t-1}\right)\|_{2}^{2}\leq\left(\mathcal{A}E-\mathcal{A}X+\mathcal{A}\sigma_{t}\right)^{2}.$
@@
$\|\mathcal{A}\left(\mathbf{x}_{t-1}^{i}-\mathbf{x}^{*}_{t-1}\right)\|_{2}^{2}$
[…]
$\overset{(\text{a})}{\approx}$ $\|\mathcal{A}\left(\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}^{i}_{t}\right)-$
(27) $\left(\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\right)\|_{2}^{2},$
@@
$\|\mathcal{A}\left(\mathbf{x}_{t-1}^{i}-\mathbf{x}^{*}_{t-1}\right)\|_{2}^{2}$
$\overset{(\text{a})}{\approx}$ $\|\mathcal{A}\left(\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}^{i}_{t}\right)-$
(28) $\left(\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\right)\|_{2}^{2}$
[…]
(29) $\left(\sigma_{t}\mathcal{A}\mathbf{z}_{t}^{i}-\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\right)\|_{2}^{2}$
[…]
(30) $\left.\|\sigma_{t}\mathcal{A}\mathbf{z}_{t}^{i}-\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\|_{2}\right)^{2}.$
Since $\|\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})\|_{2}\leq E$, $\|\bar{\alpha}_{t}\|_{2}\leq A$, $\|\bar{\alpha}_{t}\|_{2}\leq A$, $\|\bar{\alpha}_{t-1}\|_{2}\leq A,$, and $\|\mathbf{y}\|_{2}\leq Y$, we have the upper bound for $\|\mathcal{A}\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})-\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}\|_{2}$, and thus $\|\mathcal{A}\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})-\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}\|_{2}\leq\mathcal{A}E+\sqrt{A}Y$. Similarly, we have the upper bound for $\|\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\|_{2}$, which is $\|\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\|_{2}=\|\frac{\sqrt{1-\bar{\alpha}_{t-1}}\mathcal{A}\mathbf{x}_{t}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{\sqrt{\bar{\alpha}_{t}(1-\bar{\alpha}_{t-1})}y}{\sqrt{1-\bar{\alpha}_{t}}}\|\leq\mathcal{A}X+\sqrt{A}Y$. Therefore, we have:
$\|\mathcal{A}\left(\mathbf{x}_{t-1}^{i}-\mathbf{x}^{*}_{t-1}\right)\|_{2}^{2}$
$\overset{(\text{a})}{\approx}$ $\|\mathcal{A}\left(\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}^{i}_{t}\right)-$
(31) $\left(\sqrt{\bar{\alpha}_{t-1}}\mathbf{y}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathcal{A}\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{y}}{\sqrt{1-\bar{\alpha}_{t}}}\right)\|_{2}^{2}$
[…]
@@
(33) $f(\mathbf{z})=\|\mathcal{A}(\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}-C_{1}\mathbf{x}_{t})-C_{2}\mathbf{y}\|^{2}_{2},$
where $C_{1}={\sqrt{1-\bar{\alpha}_{t-1}}}/{\sqrt{1-\bar{\alpha}_{t}}}$ and $C_{2}=\sqrt{\bar{\alpha}_{t-1}}-{\sqrt{\bar{\alpha}_{t}}\sqrt{1-\bar{\alpha}_{t-1}}}/\\ {\sqrt{1-\bar{\alpha}_{t}}}$. Thus, with $M$ trials, each consisting of $N$ samples, we have the variance for $f(\mathbf{z})$, which is $\text{Var}\left(f(\mathbf{z})\right)$. We have
(34) $\text{Var}_{\text{DPS}}\left(f(\mathbf{z})\right)>\text{Var}_{\text{MC}}\left(f(\mathbf{z})\right)>\text{Var}_{\text{Ours}}\left(f(\mathbf{z})\right),$
here, $\text{Var}_{\text{DPS}}\left(f(\mathbf{z})\right)$ is the variance of DPS (Chung et al., 2023a), $\text{Var}_{\text{MC}}\left(f(\mathbf{z})\right)$ is the variance of Monte Carlo sampling, and $\text{Var}_{\text{Ours}}\left(f(\mathbf{z})\right)$ is the variance of our proximal sampling method.
@@
(35) $\hat{\mu}_{\text{MC}}=\frac{1}{N}\sum^{N}_{i=1}f(\mathbf{z}_{i}).$
@@
(36) $\hat{\mu}_{\text{Ours}}=\operatorname*{arg\,min}_{\mathbf{z}}{f(\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{N})}.$
@@
(37) $\text{Var}_{\text{DPS}}\left(f(\mathbf{z})\right)>\text{Var}_{\text{MC}}\left(f(\mathbf{z})\right)=\frac{\text{Var}_{\text{DPS}}\left(f(\mathbf{z})\right)}{N}>\text{Var}_{\text{Ours}}\left(f(\mathbf{z})\right).$
Appendix C Additional Experimental Results
Figure 8. More visual comparison of inpainting and deblurring on ImageNet.